Jim is x years old.
Mike is 2 years older than Jim.
8 years ago, Mike was twice as old as Jim.
What is Jim's age?

Found 2 solutions by jennywecas, stanbon:
Answer by jennywecas(28)

Jim=x
mike=y
x+2=y (2x)=y+8
so you make them both equal to y. to make (2x)=y+8 equal to x you subtract 8 on both sdes.now since they both equal y they equal each other so your new equation would be 2x-8=x+2 you solve for x by subtracting x from both sides and then you add 8 to both sides. x=10 so jim is 10 years old.

Answer by stanbon(75887)

Jim is x years old.
Mike is 2 years older than Jim.
8 years ago, Mike was twice as old as Jim.
What is Jim's age?
---
Mike is x+2.
---
Equation:
x+2-8 = 2(x-8)
x-6 = 2x-16
x = 10 yrs old (Jim's age now)
x-2 = 8 yrs old (Mike's age now)
====================================
Cheers,
Stan H.
============